OpenStudy (anonymous):

please refresh me on substitution method: 8x-3y=8 7x+4y+60

OpenStudy (anonymous):

8x-3y = 8 y = (8-8x)/-3 put that into either equation

OpenStudy (anonymous):

oh, I remember that. lemme see if I can work it out

OpenStudy (anonymous):

well, I've gotten it to x = 8+3y/8 and 7(8+3y/8) +4(8+3y/8) right so far?

OpenStudy (anonymous):

can someone help with my problem? please

OpenStudy (anonymous):

oops, except I just used an x in place of a y on the 4, so ignore that.